For the tax year, ABC partnership reported a 68,000 ordinary loss and a 30,000 increase in recourse liabilities for which the partners are liable. Anne a 50% partner, had an adjusted basis of 20,000 at the beginning of the year. What is Anne's allowable loss and her adjusted basis in ABC at the end of the year?

1. Allowable loss - 19,000. adjusted basis - 1,000

2. Allowable loss - 20,000. Adjusted basis - 0

3 Allowable loss - 34,000. Adjusted basis - 0

4 Allowable loss - 34,000. Adjusted basos - 1,000
Answer #1

Allowable loss = Ordinary loss * 50%

= 68,000*50% = 34,000

Adjusted basis

= [(30,000*50%)+20,000] - 34,000

= 35,000 - 34,000 = 1,000

Option 4 is the answer